# Mastering NIM and Impartial Games with Weak Neural Networks: An AlphaZero-inspired Multi-Frame Approach

### v2

@@ 6 Experiments
We evaluate the multi-frame restoration hypothesis in two complementary settings. First, we construct a supervised restoration dataset from teacher–random transitions (Section 6.2), which directly labels the restoration response. Second, we describe an AlphaZero-style teacher-guided training regime using MCTS against a noisy expert (Section 6.3). Unless otherwise stated, both one-frame (1F) and two-frame (2F) agents share the same action parameterization and network architecture (Section 6.1).
@@ 6.1.2 Network architecture
Hidden 1 Fully Connected $2Nk\to H$ LayerNorm + ReLU
@@ 6.1.3 Training targets and loss
For each training state (single-frame) or state-pair (two-frame), MCTS produces a root visit-count distribution $\pi_{\mathrm{MCTS}}$. The policy head is trained to match this distribution (cross-entropy loss). The value head is trained to predict the final game outcome from the perspective of the player to move (mean squared error on targets $\pm 1$). The overall loss is the sum of policy and value losses (optionally with standard regularization).
6.1.4 Definition of mastery
We report two complementary notions of performance. Game-level performance is measured by win rate (and Elo) under the evaluation protocols of Section 6. Move-level performance is measured by functional correctness: on a test position with nonzero nim-sum, a move is counted as correct if it produces a successor state with nim-sum zero. Positions with zero nim-sum have no winning move (Theorem 3.4); when reporting functional correctness we therefore condition on nonzero nim-sum positions.
@@ 6.2.2 Dataset generation (teacher–random–teacher)
$(p,p^{\prime})\mapsto\pi^{*}(\cdot\mid p^{\prime}),$
where $\pi^{*}(\cdot\mid p^{\prime})$ is a soft target distribution over optimal moves from $p^{\prime}$.
@@
    Define the label $\pi^{*}(\cdot\mid p^{\prime})$ as the uniform distribution over all optimal (winning) moves from $p^{\prime}$.
@@ 6.2.4 Evaluation metrics
We evaluate in three complementary ways:
@@
Given a held-out example, we decode the model’s greedy action (argmax over legal moves). We count a prediction as correct if the chosen action lies in the support of $\pi^{*}(\cdot\mid p^{\prime})$ (i.e., it is one of the optimal restoration moves). This accounts for the fact that multiple optimal moves may exist.
Gameplay strength (checkpoint tournament, greedy play).
We form a pool of 20 agents (10 checkpoints from 1F and 10 from 2F) and run a round-robin tournament. Each pair plays 10 games (total 1900 games). Games start from a uniformly random nonterminal Nim position; the starting player is randomized. Agents play greedily from their policy logits (no search). We compute Elo ratings with initial rating 1500 for each agent and update factor $K=24$.
Gameplay strength with search (MCTS-per-move tournament).
To probe how restoration knowledge interacts with search, we also evaluate the same checkpoint pool under per-move Monte Carlo tree search. On every move, the acting agent runs a fixed-budget PUCT search (200 simulations per move; exploration constant $c_{\mathrm{puct}}=1.5$) using its policy network as the prior over legal actions. Leaf values are obtained by a rollout to termination using greedy play from the two checkpoint policies (one for each side); the rollout return is backed up with standard negation across plies. We run a head-to-head cross-group tournament (10 1F checkpoints $\times$ 10 2F checkpoints, 2 games per pair; 200 games total) and compute Elo as above.
@@ 6.2.5 Results
Figure 1 shows validation restoration accuracy as a function of epoch. The 1F model remains near chance throughout training, stabilizing around 0.079 (min 0.075, max 0.081 across checkpoints on the test set). In contrast, the 2F model learns the restoration mapping and reaches essentially perfect accuracy with enough training. On the test set (Figure 2), the 2F checkpoints achieve 0.209 at epoch 1, 0.538 at epoch 23, 0.876 at epoch 45, 0.989 at epoch 67, and 0.99998 by epoch 178 (remaining $\geq 0.99993$ at epoch 200).
@@
The 2F learning curve is not monotone: it first rises to an intermediate plateau near $0.54$ and remains there for many epochs before “breaking through” to perfect restoration. This plateau is consistent with an approximate-majority shortcut on the least significant parity bit: in our generated data, the least significant bit of the post-random nim-sum is slightly imbalanced (empirically $P[\mathrm{LSB}(\mathrm{nim\_sum}(p^{\prime}))=1]\approx 0.5355$), so a policy that predicts this bit as a constant can achieve accuracy around 0.54 while still learning the easier aspects of the restoration response (e.g. which heap to modify and the higher bits of the target value). Only later does training force the model to represent the full parity structure needed for exact restoration, at which point accuracy quickly approaches 1. This behaviour reflects the broader theme of Section 5.3: approximate majority-type correlations are $\mathrm{AC}^{0}$-compatible and easy to pick up, while exact parity requires sharper global consistency.
Greedy play: two frames yield a modest Elo advantage.
In the greedy round-robin tournament (no search), 2F agents win 59.6% of cross-group games (596/1000; 95% CI 56.5%–62.6%). The mean Elo of 2F checkpoints is 1562.5 versus 1437.5 for 1F (a +125.1 Elo gap).
Search amplifies the two-frame advantage.
Under per-move MCTS with a fixed search budget (200 simulations per move), the advantage of 2F priors becomes much larger. 2F wins 79.5% of cross-group games (159/200; 95% CI 73.4%–84.5%), and the mean Elo gap increases to +175.0 (1587.5 for 2F vs. 1412.5 for 1F). Figure 3 shows the resulting Elo ordering across checkpoints.
Figure 3: Head-to-head Elo of checkpointed agents under MCTS-per-move evaluation ($N=20$, $k=4$). Each move is selected by PUCT with 200 simulations and rollouts using greedy policy play from the two checkpoints.
Overall, these results support the theoretical narrative: single-frame models fail to learn the parity-governed optimal response even under perfect supervision, while two-frame models can learn the restoration rule from transition data. Moreover, once a reliable restoration prior is learned, combining it with bounded search substantially increases effective playing strength.
6.3 Teacher-guided MCTS training against a noisy teacher
This subsection describes an AlphaZero-style training protocol in which the learner plays against a near-perfect noisy teacher and selects actions via MCTS. We include it as a controlled end-to-end baseline. In practice, bootstrapping strong play from terminal win/loss outcomes can be challenging even with a teacher; the supervised restoration benchmark of Section 6.2 provides a cleaner test of the multi-frame restoration hypothesis.
This section describes an experimental protocol designed to test the theoretical prediction of Sections 4.24.3: providing a short history window should make NIM substantially easier to learn for fixed-precision bounded-depth policy/value networks.
6.3.1 Game family and start-state distribution
We consider the family of NIM games with $N$ heaps and heap sizes in $\{0,1,\dots,2^{k}-1\}$. At the start of each training episode we sample a nonterminal position by drawing each heap size independently and uniformly from $\{0,1,\dots,2^{k}-1\}$ and resampling if all heaps are zero. For evaluation we optionally stratify results by whether the initial nim-sum is zero or nonzero (Theorem 3.4).
6.3.2 Near-perfect teacher
We use the noisy expert policy $\pi_{T}^{\varepsilon}$ from Section 5.2. When $\varepsilon=0$ the teacher plays perfectly; for $\varepsilon>0$ the teacher occasionally makes random moves, which increases the diversity of visited states and transitions.
6.3.3 Training procedure: MCTS against the teacher
We train two agents independently: a single-frame agent whose networks take the current position as input, and a two-frame agent whose networks take the previous and current positions as input. Both agents are trained in an AlphaZero-style loop:
  1. (1)
    Play episodes against the teacher. On the agent’s turn, select actions using Monte Carlo tree search (MCTS) with a fixed simulation budget.
  2. (2)
    Use the root visit-count distribution produced by MCTS as the policy training target for the corresponding state (or state-pair).
  3. (3)
[…]
  4. (4)
    Update network parameters by stochastic gradient descent on mini-batches from the replay buffer.
Rollouts and back-ups.
Within each MCTS simulation, leaf evaluation can be obtained either from the current value network or from a rollout that plays the game forward to termination using a simple default policy (e.g. the current policy network with added exploration and/or the teacher for the opponent). In all cases, the return of the simulation is backed up along the path in the standard way. No additional constraints are imposed on rollout moves beyond legality.
6.3.4 Evaluation metrics
We report:
  • Win rate vs. a perfect teacher ($\varepsilon=0$), which measures performance against optimal play.
  • Win rate vs. the noisy teacher ($\varepsilon$ matching training), which measures robustness and exploitation of occasional mistakes.
  • Head-to-head win rate and Elo difference between the 1F and 2F agents under the same search budget and alternating first player.
  • Move optimality accuracy on random test positions with nonzero nim-sum: the fraction of states where the agent’s chosen move produces nim-sum zero.
For Elo reporting we use the standard logistic model. Given an observed win rate $w$ for the two-frame agent in a head-to-head match, the implied Elo difference is
$\Delta_{\mathrm{Elo}}\,=\,-400\log_{10}\!\Big(\frac{1}{w}-1\Big),$
with $w$ clipped away from 0 and $1$ for numerical stability.
@@ 7 Discussion and broader implications
Our main message is that multi-frame representations can turn globally parity-governed decision rules into locally checkable transition rules, making them feasible for fixed-precision bounded-depth neural inference. We briefly discuss two directions where this perspective may be useful beyond the specific NIM setting.
@@ 7.1 Multi-frame approaches to mathematical discovery
The $\mathrm{AC}^{0}$ perspective is not limited to game playing. It provides a general method for diagnosing when a direct neural approach is blocked by parity-like constraints and for designing alternative representations that bypass the barrier.
This principle guided our work on discovering large Condorcet domains [12, 23, 22]. The long-standing record domain (Fishburn) involves parity-like constraints across many triples; single-frame neural search struggled even at small parameters. Multi-frame and locally checkable invariants transformed the search from an $\mathrm{AC}^{0}$-infeasible global parity constraint into tractable local constraints that guide exploration.

### v3

@@ 7 Experiments
We evaluate the multi-frame restoration hypothesis with a supervised restoration dataset from teacher–random transitions (Section 7.2) and a complementary single-frame multi-head shootout evaluation (Section 7.3). Unless otherwise stated, both one-frame (1F) and two-frame (2F) agents share the same action parameterisation and network architecture (Section 7.1).
@@ 7.1.2 Network architecture
Hidden 1 Fully Connected 1F: $Nk\to H$; 2F: $2Nk\to H$ LayerNorm + ReLU
@@ 7.1.3 Training targets and loss
For each training input (single-frame state or two-frame state-pair), the target is a soft distribution $\pi_{\mathrm{opt}}(\cdot\mid p^{\prime})$ over optimal restoration moves. The policy head is trained with cross-entropy to this target. The value head is retained for architectural consistency across agents, but the main reported restoration metric is policy correctness.
7.1.4 Evaluation criteria in this version
In this version we report two task-level notions of performance rather than full game-level Elo tournaments. Move-level restoration performance is measured by functional correctness: on a test position with nonzero nim-sum, a move is counted as correct if it produces a successor state with nim-sum zero. Positions with zero nim-sum have no winning move (Theorem 4.4); when reporting functional correctness we therefore condition on nonzero nim-sum positions. Position-level evaluation performance is measured by $P/N$ classification accuracy and $N$-position recall in FSM-controlled shootouts (Section 7.3).
@@ 7.2.2 Dataset generation (teacher–random–teacher)
$(p,p^{\prime})\mapsto\pi_{\mathrm{opt}}(\cdot\mid p^{\prime}),$
where $\pi_{\mathrm{opt}}(\cdot\mid p^{\prime})$ is a soft target distribution over optimal moves from $p^{\prime}$.
@@
    Define the label $\pi_{\mathrm{opt}}(\cdot\mid p^{\prime})$ as the uniform distribution over all optimal (winning) moves from $p^{\prime}$.
@@
Why teacher–random–teacher?
We do not sample arbitrary triples $(P,P^{\prime},a)$ uniformly at random, because the restoration claim is about a specific interaction regime: the opponent departs a $P$-position and the agent must restore it. The teacher step ensures $P$ is a genuine $P$-position reachable under optimal play, the random move injects diversity in the departure $P\to P^{\prime}$, and the final teacher move provides a correct restoration label for $P^{\prime}$.
@@ 7.2.4 Evaluation metrics
We evaluate restoration performance on held-out data:
@@
Given a held-out example, we decode the model’s greedy action (argmax over legal moves). We count a prediction as correct if the chosen action lies in the support of $\pi_{\mathrm{opt}}(\cdot\mid p^{\prime})$ (i.e., it is one of the optimal restoration moves). This accounts for the fact that multiple optimal moves may exist.
@@ 7.2.5 Results
Figure 5 shows validation restoration accuracy as a function of epoch. The 1F model remains near chance throughout training, stabilising around 0.079 (min 0.075, max 0.081 across checkpoints on the test set). Here “chance” refers to a random-legal-move baseline on the same test distribution (single-digit restoration accuracy, since only a small fraction of legal moves are restoring moves). In contrast, the 2F model learns the restoration mapping and reaches essentially perfect accuracy with enough training. On the test set (Figure 6), the 2F checkpoints achieve 0.209 at epoch 1, 0.538 at epoch 23, 0.876 at epoch 45, 0.989 at epoch 67, and 0.99998 by epoch 178 (remaining $\geq 0.99993$ at epoch 200).
@@
The 2F learning curve is not monotone: it first rises to an intermediate plateau near $0.54$ and remains there for many epochs before “breaking through” to perfect restoration. We emphasise that this is a hypothesis consistent with our data distribution; see Section 6.3 for the circuit-complexity perspective on approximate majority. This plateau is consistent with an approximate-majority shortcut on the least significant parity bit: in our generated data, the least significant bit of the post-random nim-sum is slightly imbalanced (empirically $P[\mathrm{LSB}(\operatorname{nim}(p^{\prime}))=1]\approx 0.5355$), so a policy that predicts this bit as a constant can achieve accuracy around $0.54$ while still learning the easier aspects of the restoration response. Only later does training force the model to represent the full parity structure needed for exact restoration, at which point accuracy quickly approaches 1. This behaviour reflects the broader theme of Section 6.3: approximate majority-type correlations are $\mathrm{AC}^{0}$-compatible and easy to pick up, while exact parity requires sharper global consistency.
Overall, these results support the theoretical narrative: single-frame models fail to learn the parity-governed optimal response even under perfect supervision, while two-frame models can learn the restoration rule from transition data. Moreover, once a reliable restoration prior is learned, the transition-based representation yields strong and stable restoration performance. The configuration reported here ($N=20$, $k=4$, $D=10^{6}$) is representative of a broader experimental campaign. We also ran the 1F/2F comparison across different numbers of heaps, bit-widths, dataset sizes, and hyperparameter settings (hidden width, learning rate, training epochs). The qualitative pattern is consistent throughout: single-frame models remain near chance while two-frame models learn restoration, with the plateau and breakthrough dynamics described above. Code and data for all configurations are available at https://github.com/SR123/NIM.
7.3 Single-frame evaluation via FSM-controlled shootouts (multi-head)
Purpose.
We complement the two-frame restoration study (Section 7.2) with an experiment targeting the single-frame bypass from Proposition 5.9: multiple $\mathrm{AC}^{0}$-bounded rollout heads combined with a finite-state shootout controller (Figure 3). The goal is to test the rollout interface itself (how much information can be extracted by controlled shootouts), not to claim that self-play discovers the structured heads without additional inductive bias.
Setup.
We consider NIM with $N=20$ heaps and bit-width $k=4$ (so $B=k=4$ heads), matching the encoding of Section 7.1.1. We evaluate on a balanced test set of 2000 positions with $50\%$ $P$-positions ($\sigma=0$) and $50\%$ NN-positions ($\sigma\neq 0$).
Training the rollout heads.
We train a shared-trunk network with $B=4$ policy heads by imitation learning against the structured policies $\pi_{0},\dots,\pi_{B-1}$ from Proposition 5.9. Each head $\pi_{j}$ subtracts $2^{j}$ from the leftmost heap of size at least $2^{j}$ (falling back to subtracting 1 from the leftmost nonzero heap). When multiple heaps qualify for the same phase of $\pi_{j}$ (e.g. several heaps have $h_{i}\geq 2^{j}$), any choice among them produces a rollout with the same total move count and hence the same parity outcome. We therefore train with soft targets: for each position and head $j$, the label is uniform over all equivalent actions, and the loss is cross-entropy against this soft distribution. This avoids penalising the model for an arbitrary tie-breaking choice among equally valid heaps.
The network uses the same shared-trunk architecture as Table 1 (two hidden layers of width $H=256$ with LayerNorm), with $B$ independent linear output heads each producing $N\cdot 2^{k}=320$ action logits. Training uses Adam (learning rate $10^{-3}$, batch size 1024) for 12 epochs on $2\times 10^{5}$ imitation examples.
Rollout evaluation.
Each rollout is deterministic: both players follow the same greedy policy head $\pi_{j}$ (argmax over legal actions) until termination. The rollout outcome is 1 if the first player wins (makes the last move) and 0 otherwise. Since rollouts are deterministic, a single rollout per (position, head) pair suffices.
The per-head accuracy metric is whether the model’s rollout outcome matches the teacher’s rollout outcome—the correct channel bit—not whether it matches the $P/N$ label directly. Each head $j$ extracts a specific linear functional of the nim-sum ($r_{0}=\sigma_{0}$; $r_{j}=\sigma_{j}\oplus\sigma_{0}$ for $j\geq 1$), and only the FSM combining all $B$ heads reconstructs the full $P/N$ classification.
FSM-controlled shootout evaluation.
Given a test position $P$, the evaluator runs the controller of Figure 3: it queries head $\pi_{0}$ first; if $r_{0}=1$ it immediately declares $\sigma(P)\neq 0$ (N-position). Otherwise it queries heads $\pi_{1},\pi_{2},\ldots,\pi_{B-1}$ sequentially, declaring N-position as soon as any $r_{j}=1$, or declaring P-position if all $B$ bits are 0. Early stopping reduces the average number of rollouts below $B$ on N-positions. We report (i) overall $P/N$ evaluation accuracy and (ii) mean number of shootouts used.
Results.
Table 2 reports the results.
Multi-head FSM achieves perfect classification. With all $B=4$ heads, the FSM controller correctly classifies every test position (1.000 accuracy), confirming Proposition 5.9: the $B$ structured rollout heads recover all $B$ bits of the nim-sum, and the FSM reconstructs the full $P/N$ classification. The learned model matches the teacher exactly, indicating that the soft-target imitation training produces heads whose rollout trajectories preserve the parity structure needed to extract the correct channel bits.
Single-head baseline and theoretical prediction. The single-head baseline achieves 0.768 overall accuracy and 0.536 N-position recall. These numbers have an exact theoretical explanation. With $k=4$ bits, an N-position has nim-sum $\sigma\in\{1,\dots,15\}$. Of these 15 values, exactly 8 are odd ($\sigma_{0}=1$: the values $1,3,5,7,9,11,13,15$) and 7 are even ($\sigma_{0}=0$). Head $\pi_{0}$ detects an N-position only when $\sigma_{0}=1$, so the theoretical N-position recall is $8/15\approx 0.533$ (observed: 0.536 on the finite test set). All P-positions are correctly classified (since $\sigma=0$ implies $\sigma_{0}=0$), so on a balanced test set the overall accuracy is $\tfrac{1}{2}\cdot 1+\tfrac{1}{2}\cdot\tfrac{8}{15}=\tfrac{23}{30}\approx 0.767$ (observed: 0.768). This is a quantitative instance of the single-policy bottleneck (Theorem 5.14(i)): one rollout head reveals only $\sigma_{0}$, leaving B1=3B-1=3 bits unresolved.
Connection to the two-frame restoration plateau. The $8/15$ ceiling is closely related to the $\approx 0.54$ accuracy plateau in the two-frame restoration experiment (Section 7.2, Figure 5). Both reflect the information limit of $\sigma_{0}$ alone: in the restoration setting, a model that has learned to exploit the least significant bit—but not the higher bits—can correctly restore only the fraction of positions determined by that single bit, plateauing near $0.54$. The FSM single-head recall of $8/15\approx 0.533$ is the same limitation viewed from the classification side. In both cases, the gap to perfect performance measures the missing B1=3B-1=3 higher nim-sum bits—information that the multi-head FSM recovers and the two-frame model eventually learns.
Early stopping. The average number of shootouts is 2.87 rather than the worst-case 4, because any N-position with $\sigma_{0}=1$ is caught at head 0 (one rollout); only those with $\sigma_{0}=0$ require further heads. Representativeness. The configuration reported here ($N=20$, $k=4$, $B=4$) is likewise representative. We ran the multi-head FSM evaluation for larger values of $N$ and $k$ (and correspondingly more heads) and varied hyperparameters including hidden width, learning rate, training set size, and number of epochs. The qualitative pattern is consistent: the multi-head FSM achieves perfect or near-perfect win/loss position classification whenever per-head imitation accuracy is sufficiently high, while the single-head baseline remains bounded by the $\sigma_{0}$-only ceiling. Code and data for all configurations are available at the same repository.
Method Heads Shootouts (avg.) $P/N$ Accuracy (Overall) NN-pos Recall
Single-head baseline (model) 1 1.00 0.768 0.536
Multi-head + FSM (model) 4 2.87 1.000 1.000
Single-head baseline (teacher) 1 1.00 0.768 0.536
Multi-head + FSM (teacher) 4 2.87 1.000 1.000
Table 2: Single-frame $P/N$ evaluation by FSM-controlled shootouts ($N=20$, $k=4$), reporting overall $P/N$ accuracy and $N$-position recall.
@@ 8 Discussion and broader implications
Our main message is that multi-frame representations can turn globally parity-governed decision rules into locally checkable transition rules, making them feasible for fixed-latency, fixed-scale quantised inference. We also show that a multi-policy-head architecture can break the parity barrier even in a single frame. We briefly discuss several directions where this perspective may be useful beyond the specific NIM setting.
8.1 Theoretical idealisation vs. practical architecture
The formal $\mathrm{AC}^{0}$ results describe a fixed-latency logical core with fixed-scale quantised thresholds. Our empirical networks additionally use LayerNorm and Softmax (implemented as Log-Softmax), both of which are statistical normalisation layers with global aggregation and therefore exceed strict $\mathrm{AC}^{0}$ as literal circuit operations. We treat these components as numerical stabilisation/calibration layers, distinct from the logical reasoning layers required to resolve parity. The key empirical pattern is consistent with this interpretation: in Section 7.2, the one-frame model still fails near chance despite using the same normalisation layers, while two-frame access succeeds once local transition information is exposed.
@@ 8.2 Multi-frame approaches to mathematical discovery
The $\mathrm{AC}^{0}$ perspective is not limited to game playing. It provides a general method for diagnosing when a direct neural approach is blocked by high-frequency signals (Section 6.4) and for designing alternative representations that bypass the barrier.
This principle guided our work on discovering large Condorcet domains [9, 30, 29]. The long-standing record domain (Fishburn) involves parity-like patterns and high-frequency signals across many triples; single-frame neural search struggled even at small parameters. Rather than using explicit multi-frame inputs, we used a hand-crafted AI-inspired search based on local database look-ups and locally checkable invariants. This transformed the search from an $\mathrm{AC}^{0}$-infeasible global constraint into tractable local constraints that guide exploration.
@@ 8.3 Beyond NIM: other impartial games
8.4 The multi-policy-head architecture and the single-policy bottleneck
Theorem 5.14 formalises an information bottleneck in our stylised deterministic rollout-policy interface (AlphaZero-like in spirit): with a single policy head, all rollouts under that policy reveal the same linear functional of the nim-sum (Proposition 5.10), so additional search does not recover the missing bits. A multi-policy-head architecture (Definition 5.7) provides multiple independent rollout channels; in NIM, $B$ suitably structured heads suffice to recover all $B$ bits of the nim-sum (Proposition 5.9).
This suggests a qualitative separation in search requirements: a single-policy agent must effectively brute-force missing information, whereas multiple heads can supply it directly. We leave formal training- and search-complexity bounds for future work.
More broadly, whenever optimal decisions depend on a multi-bit global invariant that is hard for $\mathrm{AC}^{0}$, adding independent policy heads is a principled architectural intervention suggested by circuit complexity.
8.5 The Optimisation Gap: Why Representation is Not Enough
Our theoretical results show that both the 2F architecture and the multi-policy-head architecture remove the representational barrier: solutions are implementable in $\mathrm{AC}^{0}$. The supervised results nevertheless indicate a complementary optimisation gap: even with the right representation, learning can pass through long intermediate plateaus before converging to exact restoration. In end-to-end self-play settings, where credit assignment is substantially noisier, this optimisation gap is expected to be even more pronounced; we therefore treat teacher-guided MCTS-style training as an important future direction rather than a completed empirical claim in this version.
While the 2F agent can represent the optimal restoration rule in $\mathrm{AC}^{0}$, discovering this rule via gradient descent from sparse rewards remains challenging. Similarly, while a multi-policy-head agent can represent the $B$ structured rollout policies, discovering them through self-play is an open problem. This distinction—between what an architecture can express and what training can find—defines the frontier of the research. We have established the necessary structural priors (history/difference access, or multiple policy heads), but future work must develop the algorithmic priors (e.g., difference-attention heads, curriculum design, auxiliary losses rewarding nim-sum-zero transitions) to make these solutions discoverable by standard RL optimisers.
8.6 Significance and impact
Our results suggest several broader implications.
  • For learning theory: sequential interaction can transform a globally hard invariant into a locally checkable transition rule, changing what weak learners can achieve. Multiple independent policy heads can extract multi-bit invariants that any single policy head provably cannot.
  • For neural architecture: short history windows (or difference features) can be a principled way to bypass parity-like barriers without scaling depth/width. When history is unavailable, multiple policy heads offer an alternative bypass within a single-frame architecture.
  • For practice: when a model fails, the cause may be structural misalignment rather than data scarcity; circuit complexity offers a diagnostic lens. The single-policy bottleneck identified here may affect AlphaZero-style systems in any domain governed by multi-bit global invariants.